Suppose that w = 2^1/2, x = 3^1/3, y = 6^1/6, and z = 8^1/8. From among these numbers, list the biggest first and the second biggest second
(1) w, x (2) x,w (3) y, z (4) x, z

Answer by Alan3354(69443)

Suppose that w = 2^1/2, x = 3^1/3, y = 6^1/6, and z = 8^1/8. From among these numbers, list the biggest first and the second biggest second
(1) w, x (2) x,w (3) y, z (4) x, z
---------------
It's x,w
---------
I did a similar problem years ago to find the max of x^(1/x).
The max is at x = e, the base of natural logs, =~2.71828
e^(1/e) =~ 1.44467 the max
3^(1/3) =~ 1.4425
2^(1/2) =~ 1.4142